Operations Report – January 08th

SOL: 10

Name of person filing report: Kasey Hilton

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Generator (hours run): 16hr 32min; Turned on last night (07Jan2019) at 16:26; Turned off this morning (08Jan2019) at 8:58; Turned on tonight (08Jan2019) at 16:33

Solar SOC – Turned on (07Jan2019) 79%; Turned off (08Jan2019) 100%; Turned on (08Jan2019) 80%

Diesel Reading – 70%

Propane Reading – 39%

Ethanol Free Gasoline – Not in use

Water (auxiliary tank) – Not in use

Water (static tank) – About 55%; 295 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 65%; 190 gallons

Water (loft) – At level marker 12

Static to Loft Pump used – Yes; At 20:01 to refill tank

Water Meter: 01398995

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: No

Hours: 66.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Opportunity rover used: No

Hours: 45.4

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: No

Hours: 68.0

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Brake fluid in Opportunity was replaced but the brakes are still not functioning properly

ATV’s Used: Honda, YODA (300), 350.1 (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: Geology EVA that was too long for rovers

Oil Added? No

ATV Fuel Used: About half a gallon; Still mostly full

# Hours the ATVs were Used today: 2.5 hrs

Notes on ATVs: ATVs returned very muddy due to the melting snow; An engineering EVA will be made to attempt to get some of the mud off of them

HabCar used and why, where? Not used

CrewCar used and why, where? Off site

General notes and comments: Nothing to report

Summary of internet: Internet usage is back to normal

Summary of suits and radios: Nothing to report

Summary of Hab operations: Side panel is no longer banging against the side of the Hab due to lack of wind; Has not been fixed yet but should be fixed in the next few days

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Hot plate is not being used until it can be cleaned

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report